# Improving Outdoor Multi-cell Fingerprinting-based Positioning via Mobile Data Augmentation

### v1

@@ IV-D Statistical Model Comparison
In order to perform the pairwise comparisons, we followed the methodology described above by performing $N_{e}=30$ independent runs for each model given a region and an augmentation rate, considering a confidence level of 95% (i.e., $\alpha=0.05$). Fig. 4 summarizes these pairwise comparisons for each region: rows and columns index augmentation rates $A\in\{1,5,10,20,30\}$, and each cell indicates whether the difference in the chosen positioning metric between the row- and column-level is statistically positive, negative, or not significant, highlighting both monotonic trends and saturation points.
The results show that in dense and small-area regions (city center and stadium scenarios) improvements are observed up to moderate augmentation (such as A=10A=10-2020) and then saturate: pairwise tests show no statistically significant gain when moving from $A=20$ to $A=30$. Similarly, in the airport scenario, the gains are marginal at high augmentation rates, underscoring that beyond a region-dependent saturation point, adding more synthetic samples yields diminishing returns and is unlikely to improve positioning further. Conversely, in the highway scenario, characterized by highly anisotropic sampling and low density, the improvement continues to larger augmentation rates and the pairwise matrices indicate significant gains up to $A=30$. Overall, these comparisons confirm that mobile data augmentation can substantially improve positioning, especially in sparse regions, and that improvements tend to plateau beyond a region-dependent augmentation level.
V Conclusion
This paper has presented a practical, modular framework for mobile data augmentation tailored to real-world MDT traces collected by MNOs. The proposed two-stage architecture decouples spatial sample generation from radio-feature synthesis through KDE and KNN, delivering a training-free, interpretable, and computationally lightweight solution that directly leverages operator-held geo-tagged measurements. Through empirical evaluations, we validated each module independently and we assessed end-to-end positioning performance across heterogeneous large-scale outdoor MDT reference scenarios. Results show consistent reductions in positioning error when using the proposed augmentation, with the largest relative gains in spatially sparse or structurally constrained regions. Competing two-stage baselines did not match these improvements at any tested augmentation rate. We also observed a clear saturation of benefits beyond a region-specific augmentation factor, which indicates that adding more synthetic samples does not by itself guarantee better accuracy and, when fingerprints are poorly matched, can even degrade it. In sum, the proposed KDE plus KNN augmentation offers a simple, interpretable, and deployment-ready path to more accurate outdoor fingerprinting on real MDT data, while providing operator-relevant guidance on when additional synthetic samples help and when they do not.

### v2

@@ VI-D Statistical Model Comparison
In order to perform the pairwise comparisons, we followed the methodology described above by performing $N_{e}=30$ independent runs for each model given a region and an augmentation rate, considering a confidence level of 95% (i.e., $\alpha=0.05$). Fig. 4 summarizes these pairwise comparisons for each region: rows and columns index augmentation rates $A\in\{1,5,10,20,30\}$, and each cell indicates whether the difference in the chosen positioning metric between the row- and column-level is statistically positive, negative, or not significant, highlighting both monotonic trends and saturation points. These results reveal a clear structural difference between the considered areas. In the city center scenario, which features the highest UE density, we observe that performance saturates starting at moderate augmentation rates ($A=10$). In denser scenarios, such as stadium and airport, the transition from $A=20$ to $A=30$ yields no significant gain, suggesting that the synthetic data has effectively filled the spatial gaps within the radio map. In contrast, the highway scenario, which features the lowest UE density, exhibits a sustained improvement, with significant gains persisting up to $A=30$. This behavior is driven by the highly sparse nature of MDT data in this area. Consequently, we conclude that the saturation point for positioning accuracy is inversely related to the initial UE density of the training data: sparse topologies (e.g., highway) exhibit sustained performance gains at higher augmentation rates, whereas dense environments rapidly converge to a saturation point.
VII Conclusion
This paper has presented a practical, modular framework for mobile data augmentation tailored to real-world MDT traces collected by MNOs. The proposed two-stage architecture decouples spatial sample generation from radio-feature synthesis through KDE and KNN, delivering an interpretable and computationally lightweight solution that directly leverages operator-held geo-tagged measurements. Extensive experimental validation on four large-scale operational datasets demonstrates that the proposed pipeline yields consistent accuracy gains, particularly in scenarios characterized by sparse sampling and high mobility. Specifically, our KDE-KNN framework reduced the median positioning error by approximately $30\%$ in the sparse highway scenario and by $23\%$ in the complex stadium environment, significantly outperforming more complex deep learning and generative baselines. Furthermore, our analysis revealed a critical structural dependency between augmentation efficacy and the native density of the environment. We observed that benefits from augmentation saturate more rapidly in dense urban centers than high-mobility and sparse regions. This underscores a key operational insight: augmentation strategies must be dynamically tuned to local data sparsity, as indiscriminate over-sampling in already dense areas yields diminishing returns and may not introduce useful information. In sum, the proposed KDE-KNN augmentation offers a simple, interpretable, and deployment-ready path to more accurate outdoor fingerprinting on real MDT data, while providing operator-relevant guidance on when additional synthetic samples effectively improve positioning.